OpenStudy (anonymous):

How is T o (S o R) = (T o S) o R? Let R be a relation from A to B S is a relation from B to C T is a relation from C to D. I basically wrote down the definition for each. T o (S o R) = {(a,d)∈AxD | ∃c∈C [ ∃b∈B ((a,b)∈R and (b,c)∈S) and (c,d)∈T] } (T o S) o R = {(a,d)∈AxD | ∃b∈B [(a,b)∈R and ∃c∈C ((b,c)∈S and (c,d)∈T] }

OpenStudy (anonymous):

besides the fact that both have (a,d)∈AxD, I'm having a bit difficulty interpreting the right side of each

OpenStudy (anonymous):

I was hesitating about removing some unnecessary parentheses because I'm afraid that might change the meaning.

OpenStudy (ikram002p):

humm r u sure u need to prove ? or ur not sure if the statment is true ?

OpenStudy (anonymous):

well, yes, i'm basically trying to prove the theorem. Sadly it wasn't proved in the book and left as an exercise -.- Typical in a text book haha

OpenStudy (ikram002p):

humm what i really see is giving counter example ! cuz i dnt thnk its true sentence :O what do u think ?

OpenStudy (kinggeorge):

Composition of functions is always associative. But defining them in terms of relations has always been a bit confusing for me. And I've never seen a proof of this fact using that specific notation.

OpenStudy (anonymous):

it is a theorem (as stated in the book). So T o (S o R) = (T o S) o R is true. I just can't to make sense of the definition I wrote

OpenStudy (ikram002p):

intresting ... but i cant focus with music :D https://www.youtube.com/watch?v=rYEDA3JcQqw

OpenStudy (anonymous):

haha I'm watching people playing video games while proving the theorem XD

OpenStudy (kinggeorge):

Looking at the definitions you've written down though, I'm relatively positive that you can get rid of some of the parentheses and re-order some of the statements that don't depend on one another. For example, "∃b∈B [(a,b)∈R and ∃c∈C ((b,c)∈S and (c,d)∈T]" could be rewritten as "∃b∈B and ∃c∈C [(a,b)∈R ,(b,c)∈S and (c,d)∈T]"

OpenStudy (anonymous):

@KingGeorge yes, I thought that removing some parenthesis and changing the order of each statement within the set should not change the meaning. But then again, when I have doubts, I don't wanna to make an invalid conclusion

OpenStudy (kinggeorge):

And from there you can switch the order of "∃b∈B and ∃c∈C" to "∃c∈C and ∃b∈B" and do similar rearranging of "(a,b)∈R ,(b,c)∈S and (c,d)∈T." You can make these rearrangements because the only things that might cause trouble, are moving quantifiers (∃ symbols) around statements such as "(a,b)∈R." The only time I did this, was moving a quantifier only involving the element "c∈C" around a statement not involving that same element. So in that case, we can switch the order.

OpenStudy (kinggeorge):

So doing this, you can rewrite the set-theoretic definition of (T o S) o R to be exactly the same as the set-theoretic definition of T o (S o R). Since the two sets are then equal, the theorem is proved.

OpenStudy (anonymous):

As far as I know, the order of the qualifies, if all of them are the same, does not matter. However, if there mix then it really changes the meaning. For example: ∃x∀y is completely different from ∀y∃x

OpenStudy (kinggeorge):

That's also true. But we only have one type of quantifier here, so we don't have to worry about that.

OpenStudy (anonymous):

I might be thinking too hard. But as you know, in mathematics, sometimes things that we thought (for sure) that they're true can turn out to be false and counter intuitive after proving them. I guess I'll just take their words (the mathematicians') and believe that the theorem is true hehe

OpenStudy (kinggeorge):

I think my favorite example of how important these quantifiers are, is the following question. Is this statement true or not?\[\forall w\in\mathbb{R},\exists x\in\mathbb{R},\forall y\in\mathbb{R},\exists z\in\mathbb{R}(wxyz=w+x+y+z).\]What about\[\exists w\in\mathbb{R},\forall x\in\mathbb{R},\exists y\in\mathbb{R},\forall z\in\mathbb{R}(wxyz=w+x+y+z)?\]

OpenStudy (anonymous):

O.O torture!!!!! XD

OpenStudy (anonymous):

I can't even begin to think XD

OpenStudy (the_fizicx99):

I'm not exactly here yet >.> but wouldn't wxyx = w * x*y*x?

OpenStudy (anonymous):

This is why sometimes I hate mathematics. it's only when you really understand something 100% can you be confident that it's correct. My motto is that, if there is even a slightest doubt, it means I still don't understand it completely.

OpenStudy (ikram002p):

well , sourwing i love math for that reson only :D understanding concepts is much more imp than giving a solution :D

OpenStudy (kinggeorge):

While that may be true, there comes a point when you just have to move on and accept that some things you just don't understand completely, and there's nothing wrong with that.

OpenStudy (anonymous):

hehe, which is why I'm about to turn on my PS3 and starting trolling with some noobs in Call of duty ahaha!! XD

OpenStudy (anonymous):

anyways, thank you for your time @KingGeorge @tHe_FiZiCx99 @ikram002p

OpenStudy (the_fizicx99):

CoD >.>

OpenStudy (kinggeorge):

You're welcome. And just btw, the first statement I put up there is true, while the second is not.

OpenStudy (ikram002p):

no lol i dint help , i enjoyed with the problem :D

OpenStudy (the_fizicx99):

Pardon my ignorance but how is w + x + y + z = wxyz?

OpenStudy (kinggeorge):

In general, you're correct. But by choosing specific values for \(x\) and \(z\), I can make it true.

OpenStudy (the_fizicx99):

I thought about that as well, but it seemed a bit more work. I don't know how to explain it though :/ I'm horrible with words.

OpenStudy (kinggeorge):

All you have to do, is choose \(x=0\), and \(z=-w-y\). Then both sides are 0. And since \(w\) and \(y\) have already been chosen, \(-w-y\) is a specific value.

OpenStudy (ikram002p):

check this intrested vedio. sourwing hope u dnt mind posting this in ur thread https://www.youtube.com/watch?v=w-I6XTVZXww

OpenStudy (kinggeorge):

In the second statement though, we can't do that process anymore.

OpenStudy (the_fizicx99):

Seems interesting. ^_^